Beginning Inventory Jan. 1 (carried over from 2017) 20 units @ $10

Purchase Jan 25 2018 40 units @ 11

Purchase June 20 40 units @ 12

Purchase Oct 10 50 units @ 13

Assuming the perpetual moving-average cost flow assumption, what is the company's cost of goods sold for the 120 items sold in 2018? a. $1,386 b. $1,410 c. $1,416 d.$1,460
Answer #1

Cost of Goods available for sale = (20*10)+(40*11)+(40*12)+(50*13)

= 1770

Units available = 20+40+40+50 = 150

Average cost per unit = 1770/150 = 11.80

Cost of Goods Sold = 11.80*120

= 1416

Option C